OpenStudy (anonymous):

PLEASE HELP!!!!!! For research purposes, a tornado can often be modeled as a right circular cylinder that rotates about its axis. A cylindrically shaped tornado has a core diameter of 200 ft. If the tornado has a maximum wind speed of 300 ft per second at the outer edge of the core, approximate the number of revolutions the core makes each minute. Approximate to the nearst unit

OpenStudy (anonymous):

So 300 ft per secon is the speed at the outer edge. Like this

OpenStudy (anonymous):

You have to find out how long does it take to do a full turn To do so, you must know the circumference, which is \[2 \cdot \pi \cdot r = 2 \cdot \pi \cdot 100 = 200\pi\] If it travels at 300 ft per second, it can make \[200\pi\] in \[200\pi\over 300\] seconds.which is aproximately 2 seconds. If it takes it 2 seconds to do a full turn, than it rotates 30 times a minute